OpenStudy (anonymous):

What is the equation of the line that is perpendicular to -x + y = 7 and passes through (-1, -1)

OpenStudy (anonymous):

to do this, the first thing you need is the slope of \(-x+y=7\) do you know it?

OpenStudy (anonymous):

oh.. no..

OpenStudy (anonymous):

ok then lets go slow so you get all the steps

OpenStudy (anonymous):

to find the slope of \(-x+y=7\), solve the equation for \(y\) in one step by adding \(x\) to both sides, giving \[y=x+7\]

OpenStudy (anonymous):

\[y=x+7\] looks like \[y=mx+b\] where \(m\) is the slope. you don't actually see a number in front of the \(x\), this makes \(m=1\) i.e. the slope is \(1\) ok so far?

OpenStudy (anonymous):

yeah..

OpenStudy (anonymous):

then you want the slope of the perpendicular line that is the "negative reciprocal" of the slope of this line the negative reciprocal of \(1\) is \(-1\) so the slope of the line you are looking for is \(-1\)

OpenStudy (anonymous):

the slope of the line is \(-1\) and the point is \( (-1, -1)\) use the "point slope" formula and get \[y-(-1)=-1(x-(-1))\] more succinctly written as \[y+1=-1(x+1)\]

OpenStudy (anonymous):

but its giving me multiple choice answers and thats not one of them..

OpenStudy (anonymous):

probably not that is just what you get when you write the point slope formula you can solve this equation for \(y\) if that is what you need